OpenStudy (anonymous):

A water tank is in the form of a right circular cylinder with height 20 ft and radius 6 ft. If the tank is half full of water, find the work required to pump all of it over the top rim. For this one I am slightly confused. I know it should be something like: int(0,20) 62.4 * 36pi * (10-y) dy That gives me 0 though. I've tried integrating from 10 to 20 also. I can't figure it out, the answer is 1058591.1 ft-lbs

OpenStudy (amistre64):

how much does water weigh?

OpenStudy (amistre64):

a right circular cylindar is a tin can; the crossection at any given moment is just the area of the circle

OpenStudy (anonymous):

62.4 lb/ft is the weight of water the area of one slice of the cylinder is pi*r^2 which is 36pi

OpenStudy (amistre64):



OpenStudy (amistre64):

Work = Force times distance = (2pi r)(h)(weight)

OpenStudy (amistre64):

2 pi 6 = 12pi, not 36 pi

OpenStudy (anonymous):

why not 36pi? I need pi*r^2 * dy for the volume of a slice of water. If I did 2 pi 6, wouldn't that give me the circumference?

OpenStudy (amistre64):

pfft, yeah youve got the better idea :)

OpenStudy (anonymous):

the height is the part I can't figure out. It's not (10-y) because that makes the integral 0. \[\int\limits_{0}^{20}62.4\times36\pi(height)\]

OpenStudy (amistre64):

area * height pi r^2 dh; 36pi dh from 10 to 20, since it is half full

OpenStudy (anonymous):

whoops forgot dy

OpenStudy (phi):

It looks like each slice should be raised 20-y where y is the height of the slice integrate from 0 to 10 (the highest slice)

OpenStudy (amistre64):

\[64.4(36pi)\int_{10}^{20}h~dh\]

OpenStudy (anonymous):

ah @phi yes! no wonder. thank you for your help too @amistre64

OpenStudy (amistre64):

\[62.4(36pi)(\frac{20^2-10^2}{2})\] \[62.4(18pi)(20^2-10^2)\] \[62.4(18pi)(300)=1,058,591.06\]

OpenStudy (anonymous):

I'll be posting another question rather similar to this in just a second, for clarification.

OpenStudy (phi):

with a change of variables (and limits) you get amistre's set up